OpenStudy (vshiroky):

Find the equation of the line that is perpendicular to x+2y=-9 and that passes through (-3,-4). Express your answer in standard form. What is the standard form of the equation of the line that satisfies the given properties?

jimthompson5910 (jim_thompson5910):

able to get anywhere?

OpenStudy (vshiroky):

I'm pretty lost I just changed numbers

jimthompson5910 (jim_thompson5910):

First you need to solve x+2y=-9 for y

jimthompson5910 (jim_thompson5910):

x+2y=-9 2y=-9-x 2y = -x - 9 y = -1/2x - 9/2 The slope of that line is ???

OpenStudy (vshiroky):

2

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

y = -1/2x - 9/2 is in the form y = mx+b where m = -1/2 b = -9/2

OpenStudy (vshiroky):

its a perpendicular line so my lecture said I have to flip the -1/2

jimthompson5910 (jim_thompson5910):

oh i see what you are saying (I was asking about the initial line)

jimthompson5910 (jim_thompson5910):

yes, the perpendicular slope is 2/1 or 2

jimthompson5910 (jim_thompson5910):

y = mx + b y = 2x + b .... plug in that perpendicular slope -4 = 2(-3) + b ... plug in (x,y) = (-3,-4) now solve for b

jimthompson5910 (jim_thompson5910):

What do you get when you do so?